in year 2014, rob's age was 30 years old while kobe is 23 years old .in what year when rob was three halves as old as kobe?

Answer by stanbon(75887)

in year 2014, rob's age was 30 years old while kobe is 23 years old .in what year will rob be three halves as old as kobe?
-----
Note: Rob is 30-23 = 7 yrs older than Kobe.
------
Equation:
Let the number of years be "x"::
30+x = (3/2)(23+x)
----
60+2x = 69 + 3x
----
x = -9
-----
Ans: 2014 - 9 = 1995 (the year rob will be (3/2)Kobe's age
-----
Cheers,
Stan H.
------